How would you prepare 200 mL of 0.1 M potassium acetate buffer at pH 4.5 (please give with clear detailed steps)
Answer #1

Concentration of buffer = 0.1 M

[acetic acid] + [acetate] = Concentration of buffer

[acetic acid] + [acetate] = 0.1 M    ...(1)

pKa of acetic acid = 4.76

According to Henderson - Hasselbalch equation,

pH = pKa + log([conjugate base] / [weak acid])

pH = pKa + log([acetate] / [acetic acid])

4.5 = 4.76 + log([acetate] / [acetic acid])

log([acetate] / [acetic acid]) = 4.5 - 4.76

log([acetate] / [acetic acid]) = -0.26

[acetate] / [acetic acid] = 10-0.26

[acetate] / [acetic acid] = 0.55   ...(2)

Solving equations (1) and (2) simultaneously,

[acetic acid] = 0.0645 M

[acetate] = 0.0355 M

moles acetic acid = (molarity acetic acid) * (volume of buffer in Liter)

moles acetic acid = (0.0645 M) * (0.400 L)

moles acetic acid = 0.0258 mol

mass acetic acid = (moles acetic acid) * (molar mass acetic acid)

mass acetic acid = (0.0258 mol) * (60.05 g/mol)

mass acetic acid = 1.55 g

moles acetate = (molarity acetate) * (volume of buffer in Liter)

moles acetate = (0.0355 M) * (0.400 L)

moles acetate = 0.0142 mol

mass potassium acetate = (moles acetate) * (molar mass potassium acetate)

mass potassium acetate = (0.0142 mol) * (98.14 g/mol)

mass potassium acetate = 1.39 g

To prepare the buffer solution, mix 1.55 grams of acetic acid and 1.39 grams of potassium acetate in 400 mL water.

moles acetate =